OpenStudy (anonymous):

What is the equation of the line that passes through the point (1, 6) and has a slope of 3?

OpenStudy (anonymous):

y = 3x + 5 y = 3x + 3 y = 3x - 17 y = 3x - 5

OpenStudy (anonymous):

Use the equation y=mx+b and plug in the slope, m, and the points x and y to solve for b.

OpenStudy (anonymous):

how do i set up mx+b when i don't have a second number?

OpenStudy (anonymous):

so you know Y=6 X=1 and M=3 so then you plug those in. 6=3(1)+b now you are able to solve for b. do that and then tell me what you get for b

OpenStudy (anonymous):

then we will go onto the final step

OpenStudy (anonymous):

were you able to find b?

OpenStudy (anonymous):

not really :(

OpenStudy (anonymous):

It's find we will work through it together.

OpenStudy (anonymous):

So, you have the equation: 6=3(1)+b and now you need to simplify both sides of the equation which would give you 6=3+b now you need to isolate b and you do that by subtracting 3 from both sides which would give you 3=b now you have solved for b The final step is to plug b and m back into the equation giving you the final answer of y=3x+3

OpenStudy (anonymous):

Thanks! I understand more now. Can you please help me with another?

OpenStudy (anonymous):

Sure! what is your question?

OpenStudy (anonymous):

its kind of similar: What is the equation of the line that passes through the points (-3, 2) and (-5, 8)?

OpenStudy (anonymous):

what we would do now is line them up in mx+b format

OpenStudy (anonymous):

For this problem, I would use the Point-Slope form. Are you familiar with this form?

OpenStudy (anonymous):

Somewhat

OpenStudy (anonymous):

Okay. so we will start with the equation. (first y- second y)=m(first x- second x) because we now lets plug in the variables we already have.

OpenStudy (anonymous):

(2-8)=m(-3--5) which will give us -6=m(2) or -6=2m from there we want to isolate m so we divide both sides by 2 giving us -3=m now we have solved for m

OpenStudy (anonymous):

Do you need the equation in y=mx+b form or point slope form?

OpenStudy (anonymous):

Is there a right or wrong answer to that? :s

OpenStudy (anonymous):

Nope! I was just wondering if your class required it to be in a certain form. I can show you both ways!

OpenStudy (anonymous):

There's no specific way for it to be done. I would prefer mx+b because I'm most familiar with that, but then again it would be best to learn other ways too

OpenStudy (anonymous):

So, now that you have all of the variables, you can plug them into the equation. (you can chose which coordinate pair to plug in because they will both give you the same answer) So point slope form would look like this: (y-2)=-3(x--3) or (y-2)=-3(x+3) now once you have it in this form you are able to change it into y= mx+b form

OpenStudy (anonymous):

Here's how

OpenStudy (anonymous):

First you need to simplify both sides of the equation so you will end up with y-2=-3x-9 next you need to isolate y so you add to both sides which would give you y=-3x-7 and that is your final answer now that it is in y=mx+b form!

OpenStudy (anonymous):

i get what you're saying, but that's not a choice for the question

OpenStudy (anonymous):

oh. maybe i made a math mistake. what are your options?

OpenStudy (anonymous):

3x - y = -11 3x + y = - 7 3x + y = 5 x + 3y = 3

OpenStudy (anonymous):

B is similar

OpenStudy (anonymous):

the answer i got was correct and you are right it is b. The reason they look different is because the options are in standard form (ax+by=c) so if you solved for c, which in this case is -7 you would get the answer 3x+y=-7

OpenStudy (anonymous):

That's what I figured :) thanks so much for helping me! before you go, would I solve this in the same way? Select the equations that are parallel and perpendicular to y = -3x - 1 and that pass through the point (3, 1).

OpenStudy (anonymous):

Yes! it would be similar to the first problem we solved. But first you have to find what the slopes would be for each line. Do you know how to do that?

OpenStudy (anonymous):

maybe -3 - 3 -------- 1 - 1

OpenStudy (anonymous):

so yes that is an option for finding the slope. but because you know the lines are going to be parallel and perpendicular there is an even easier way to figure them out. If lines are parallel, then they have the same slope. So, for the parallel line the slope will be -3. When lines are perpendicular, the slope will be the opposite and the reciprocal. So for the perpendicular line, the slope would be 

OpenStudy (anonymous):

now that you have the slopes, you can find both of the lines just like in the first problem!

OpenStudy (anonymous):

parallel: y = -3x + 6 perpendicular: y = 1/3 x + 2 2/3 ?

OpenStudy (anonymous):

is this correct?

OpenStudy (anonymous):

would you mind typing out the equations you used because i'm getting different answers

OpenStudy (anonymous):

i'm not really sure, i tried to solve this mentally

OpenStudy (anonymous):

did you get options for this problem?

OpenStudy (anonymous):

yes